OpenStudy (anonymous):

Prove: Suppose that f(x)=g(x)q(x)+r(x). Then d(x) is a gcd of f(x) and g(x) iff d(x) is the gcd of g(x) and r(x).

OpenStudy (anonymous):

let gcd of [g(x) and r(x)]=z(x) so g(x)=z(x)l(x) r(x)=z(x)k(x) f(x)=g(x)q(x)+r(x) f(x)=z(x)[l(x)q(x)+r(x) so gcd of f(x) and g(x) is z(x)